OpenStudy (anonymous):

Can someone help me Write 0.000032 in scientific notation?

OpenStudy (ranga):

\[\Large 3.2 \times 10^{-5}\]

OpenStudy (ranga):

In scientific notation, the first number should be between 1 and 10 followed by X followed by 10 raised to some exponent. 0.000032 = 0.000032 / 1 Multiply top and bottom by 10^5 3.2 / 10^5 = 3.2 X 10^(-5)

OpenStudy (anonymous):

Thanks

OpenStudy (ranga):

You are welcome.